OpenStudy (anonymous):

f(x) =ln(x^2 + 4) f'(x) = 2 / x^2 +4 f'' (x) = 2(4 - x^2) / (x^2 + 4) What are critical numbers?

OpenStudy (anonymous):

\[f \prime \left( x \right)=\frac{ 2x }{ x ^{2} +4}\]

OpenStudy (anonymous):

\[f \prime \prime \left( x \right)=\frac{\left( x ^{2}+4 \right)2-2x*2x }{ \left( x ^{2}+4 \right)^{2} }\] \[f \prime \prime \left( x \right)=\frac{ -2x ^{2}+8 }{\left( x ^{2}+4 \right) ^{2}}\]